OpenStudy (anonymous):

In the expansion of (a+x)^6 the first three terms are 1+3+3.75. Find the values of a and x.

OpenStudy (noelgreco):

You must have been studying binomial expansions: that will give you the coefficients for terms 2 and 3.

OpenStudy (anonymous):

1=6C0(a^6)? 1=a^6 a=1?

OpenStudy (anonymous):

yup i got it

OpenStudy (paxpolaris):

good

OpenStudy (anonymous):

and then 3=6C1(a^6-1)(x^1) 3=6x x=0.5

OpenStudy (anonymous):

awesome i got it.

OpenStudy (paxpolaris):

correct